OpenStudy (crashonce):

The equation x^2 = (y+1)^2 - y^2, where x and y are both positive integers has an infinite number of solutions. (i) One solution is x=3, y=4. Find six other solutions. (ii) Write down six solutions to the equation x^2 + y^2 = z^2, where x, y and z are all positive integers with no factors in common and where x

OpenStudy (crashonce):

@zzr0ck3r

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\color{red}{1.)}\hspace{.33em}\\~\\ x^2=(y+1)^2-y^2\hspace{.33em}\\~\\ x^2=y^2+2y+1-y^2\hspace{.33em}\\~\\ x^2=2y+1\hspace{.33em}\\~\\ \dfrac{x^2-1}{2}=y\hspace{.33em}\\~\\ \dfrac{(x-1)(x+1)}{2}=y\hspace{.33em}\\~\\ \normalsize \text{here if u choose even integers,}\hspace{.33em}\\~\\ \normalsize \text{ there will be no integers solution for y}\hspace{.33em}\\~\\ \normalsize \text{so choose only odd integers above 3, }\hspace{.33em}\\~\\ \normalsize \text{as 3 is already given }\hspace{.33em}\\~\\ \normalsize \text{example }\hspace{.33em}\\~\\ \dfrac{(5-1)(5+1)}{2}=y\hspace{.33em}\\~\\ y=12 }\end{align}\)

ganeshie8 (ganeshie8):

looks neat xD that also proves that every odd integer can be expressed as difference of two squares

OpenStudy (mathmath333):

thnks

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ \color{red}{2.)}\hspace{.33em}\\~\\ \normalsize \text{for 2 }\hspace{.33em}\\~\\ \normalsize \text{we can use the same solution as 1 }\hspace{.33em}\\~\\ x^2 = (y+1)^2 - y^2 \hspace{.33em}\\~\\ x^2+y^2=(y+1)^2\hspace{.33em}\\~\\ \normalsize \text{put z=x+1 }\hspace{.33em}\\~\\ \normalsize \text{use the same process as stated in }~~\color{red}{1.)}\hspace{.33em}\\~\\ \normalsize \text{example,for x=5 }\hspace{.33em}\\~\\ \dfrac{(x-1)(x+1)}{2}=y\hspace{.33em}\\~\\ \dfrac{(5-1)(5+1)}{2}=y\hspace{.33em}\\~\\ y=12 \hspace{.33em}\\~\\ 5^2+12^2=13^2\hspace{.33em}\\~\\ \gcd (5,12,13)=1\hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (michele_laino):

Please note that, another way to answer to the question i) is this: I call x with m, and y with n, so I can write: \[m ^{2}=(n+1)^{2}-n ^{2}=2n+1\] So we have to find all perfect square and odd numbers. For example, we have: \[9 \; \rightarrow m=3,\quad n=4\] \[25 \; \rightarrow m=5,\quad n=12\] \[49 \; \rightarrow m=7,\quad n=24\] \[81 \; \rightarrow m=9,\quad n=40\] \[121 \; \rightarrow m=11,\quad n=60\] \[169 \; \rightarrow m=13,\quad n=84\] \[225 \; \rightarrow m=15,\quad n=112\] Since there are infinite odd numbers, then there are infinite solutions.